OpenStudy (anonymous):

HELP!!!!!! What is the value of the 11th term in the sequence -3, -6, -12, -24, ...?

OpenStudy (anonymous):

first ask yourself what is the pattern?

OpenStudy (anonymous):

it looks to me that every time the value is multiplied by 2 -3X2=-6 -6X2=-12 -12X2=-24 so now you just -3 x 2^n

OpenStudy (anonymous):

use n=11 to find your solution or just continue to multiply by 2 until you arrive at the 11th number in the sequence

OpenStudy (anonymous):

Thank You!!!!